# An Efficient and Adaptive Watermark Detection System with Tile-based Error Correction

### v1

@@ Abstract
Efficient and reliable detection of generated images is critical for the responsible deployment of generative models. Existing approaches primarily focus on improving detection accuracy and robustness under various image transformations and adversarial manipulations, yet they largely overlook the efficiency challenges of watermark detection across large-scale image collections. To address this gap, we propose QRMark, an efficient and adaptive end-to-end method for detecting embedded image watermarks. The core idea of QRMark is to combine QR Code–inspired error correction with tailored tiling techniques to improve detection efficiency while preserving accuracy and robustness. At the algorithmic level, QRMark employs a Reed–Solomon error correction mechanism to mitigate the accuracy degradation introduced by tiling. At the system level, QRMark implements a resource-aware stream allocation policy that adaptively assigns more streams to GPU-intensive stages of the detection pipeline. It further employs a tile-based workload interleaving strategy to overlap data-loading overhead with computation and schedules kernels across stages to maximize efficiency. End-to-end evaluations show that QRMark achieves an average $2.43\times$ inference speedup over the sequential baseline.
@@ 1 Introduction
Recent advances in diffusion-based generative models [31, 10, 25, 32] have progressively narrowed the gap between synthetic outputs and real images, making them increasingly difficult to distinguish. In this context, watermarking has become a crucial tool for verifying AI-generated content and supporting copyright attribution. Today, major social platforms handle billions of new visual uploads each day. For example, YouTube reports over 500 hours of video uploaded per minute on its official site [44], and a recent measurement study estimates approximately 117 million TikTok posts per day [33]. Historical records also show Facebook receiving more than 350 million new photos per day [38]. The growing scale of online content creation highlights the need for watermark detection pipelines that not only achieve high accuracy, but also maintain robustness against adversarial perturbations and provide high efficiency when processing massive volumes of images.
@@
Among existing watermarking techniques, in-generation watermarking [9, 41, 39], where the watermark is incorporated directly into the generative process, has received the most attention and has emerged as the most widely adopted solution due to its accuracy and robustness. However, existing efforts largely overlook two critical practical factors of this method: the latency of watermark detection and the throughput required to reliably handle the massive volume of incoming images at scale. For example, even for relatively small $256\times 256$ images, the representative Stable Signature model requires about $14$ minutes to process $1$ million images—far from sufficient for social platforms that must process continuous streams of large-scale image and video content.
@@
One potential strategy to reduce computational overhead is tiling, where the detection process operates on a tile that still contains the embedded message. While tiling can substantially lower the computational cost of watermark detection, naive tiling does not necessarily yield proportional performance gains and often results in a noticeable drop in accuracy and robustness. Our experiments show that using the original pretrained Stable Signature decoder for tile-level detection achieves only $87.5\%$ bit accuracy, compared to $99.7\%$ with the full Stable Signature pipeline. Additional experiments consistently confirm that tiling inevitably degrades detection performance, underscoring the intrinsic trade-off between computational efficiency and accuracy. In this context, we require a method that combines low computational overhead with strong error-correction capability to offset the accuracy loss induced by tile-based watermarking. A familiar example is the QR Code, which remains reliably decodable even under partial occlusion or physical damage, thereby ensuring preservation of the stored data. Motivated by this observation, we examined the underlying mechanism and found that this resilience is largely enabled by Reed–Solomon (RS) error correction [29, 18]. We observe that RS correction holds potential for mitigating the accuracy and robustness degradation caused by tiling, as it can flexibly recover the original message bit string from partially corrupted outputs.
@@
Nevertheless, fully harnessing this potential introduces several system-level challenges. Our experiments show that applying a straightforward tiling strategy to Stable Signature yields modest performance gains ($1.17\times$ speedup) but still falls short of expectations. Moreover, RS correction is executed on the CPU—a common design choice since the algorithm is traditionally CPU-bound due to its many interdependent instruction flows (e.g., if–else branches)—and thus benefits little from GPU acceleration, which relies on massively parallel execution. In a basic pipeline, this CPU stage can easily become a bottleneck, as it must synchronize with the rest of the detection process running on the GPU. These observations underscore the need to carefully redesign the pipeline to reduce computational overhead and ensure scalability under high-throughput workloads.
@@
Beyond algorithmic design, QRMark introduces several system-level innovations to maximize GPU utilization and reduce end-to-end latency. First, we implement an adaptive CUDA stream allocation strategy that dynamically adjusts to image resolution and message length, ensuring balanced parallelism without incurring excessive kernel-launch overhead. Second, we design a resource-aware scheduling algorithm for tile-level workloads that distributes computation across GPU streams to mitigate imbalance and improve throughput. Finally, QRMark implements workload interleaving to overlap data loading with GPU execution, alleviating pipeline stall bubbles and enabling more efficient kernel orchestration.
@@
  • We design an end-to-end tile-based watermark detection framework that builds on this method while also incorporating system-level optimizations, including adaptive CUDA stream allocation and resource-aware scheduling (§5§6), to effectively overcome the inefficiencies inherent in naive tile-based detection pipelines.
@@ 2 Background
Diffusion Watermarking: Diffusion generative models [7, 10, 25, 32] have renewed interest in watermarking as a means of ensuring content provenance. In this context, watermarking refers to embedding hidden identifiers into AI-generated images, enabling origin verification without relying on visible tags. While traditional methods embed secret pixel patterns [4, 37], modern approaches integrate watermarking directly into the generative process to improve robustness. Existing image watermarking techniques can be broadly classified into three major categories: post-generation, pre-generation, and in-generation. (1) Post-generation watermarking, where the watermark is embedded after the image has been generated [28, 48], is straightforward to implement but highly vulnerable in white-box scenarios, as the watermark can be easily removed. (2) Pre-generation watermarking, which embeds watermarks into all training images [46, 50], incurs prohibitively high computational costs for large-scale diffusion models. (3) In-generation watermarking, where the watermark is embedded during the image generation process [28, 35, 52, 22, 3, 14, 47], distributes the signal across the entire image, making it difficult to remove without causing noticeable degradation. We adopt in-generation watermarking as the focus of our study, since it is widely accepted and provides superior robustness compared to alternative approaches.
System Support for Diffusion Watermarking: Prior system research on diffusion generation has primarily focused on accelerating the long sequential denoising pipeline, introducing optimizations such as kernel fusion, pipeline parallelism, stream batching, and patch parallelism [26, 17, 16, 51]. These methods improve forward sampling efficiency but assume a monolithic, step-driven workload with strong inter-step dependencies, leaving decode-time scheduling demands unaddressed. Interleaving across CPU and GPU stages is crucial for diffusion watermark detection to achieve lower latency, yet this aspect is absent from prior diffusion system designs. System-level efforts in watermarking, on the other hand, have largely targeted post-hoc embedding and detection, typically modeled as simple feed-forward computations with regular GPU kernels and minimal CPU involvement [12, 49, 36]. These methods assume image-level operators and homogeneous kernels, making them well-suited for global transforms but incompatible with patch-based workloads. In contrast, diffusion watermark detection relies on patch extraction, fragmenting computation into numerous small kernels and inducing irregular device–host interactions. Therefore, diffusion watermark detection presents new opportunities when treated as an integrated workload rather than two disjoint problems.
@@ 3 Motivation
Existing watermarking systems [9, 41, 8, 39, 50, 23] largely overlook efficiency considerations, leaving substantial opportunities for further computational optimization. To better understand how watermarking systems can be made more efficient, we focus on a representative model, Stable Signature [9], and systematically explore both algorithmic and system-level optimizations carefully built on top of it.
Divisible image watermark detection with tiling: By analyzing the Stable Signature watermark detection pipeline, we identify an opportunity to exploit tiling. Since an image is inherently divisible, it can be partitioned into tiles for localized processing. Empirical evidence from Stable Signature shows that under a cropping attack where only $10\%$ of the original area is preserved, the model still achieves over $95\%$ detection accuracy [9]. This observation suggests that, even without explicit tile-level training, the message extractor is capable of recovering the embedded message from partial image regions. Leveraging this property, we introduce tiling into the detection process. Tiling is a well-established technique in system optimization [43, 30, 13, 27], widely used to reduce both computational and memory overhead. In the context of image watermark detection, our experiment (Fig. 2(a)) demonstrates that integrating tiling into Stable Signature yields notable scalability benefits: with tiling, the system achieves up to a $1.23\times$ speedup over the non-tiled baseline, whereas the non-tiled baseline achieves at most $1.06\times$ across varying batch sizes. This result highlights the superior scalability brought by tiling. However, tiling inevitably introduces an efficiency–accuracy trade-off, as partitioning the image may compromise detection fidelity. To address this challenge, we draw inspiration from the design of QR Codes, where Reed–Solomon error correction ensures reliable decoding under data loss and corruption. This motivates our design of QRMark, which integrates Reed–Solomon correction to offset the accuracy degradation caused by tiling and achieve a more favorable balance between efficiency and robustness.
@@
Performance dilemma of watermark image tiling: Although replacing full-image processing with tile-level extraction is expected to reduce per-image overhead and improve detection efficiency, our experiments on 1,000 images from the MS-COCO [19] dataset confirm that the naive tile-based design brings only limited improvements: with a tile size of 64 and a batch size of 16, the pipeline achieves a $1.17\times$ speedup over the non-tiled baseline. However, this gain falls short of expectations, prompting the question of what factors constrain the potential of the tile-based approach. To answer this, we conducted profiling experiments with a default tile size of 64, a message length of 48, and a batch size of 16. The results reveal two primary bottlenecks. First, GPU utilization remains low: Stable Signature leverages only $31.6\%$ of available GPU capacity. Second, the Reed–Solomon correction, executed on the CPU, introduces additional overhead and occasionally stalls the entire pipeline. Further experiments also show that the transform phase preceding detection in the original Stable Signature implementation launches a large number of fine-grained kernels. These fragmented kernels originate from preprocessing operations such as cropping, resizing, and normalization. The fragmented design incurs substantial kernel-launch overhead and prevents the GPU from achieving high occupancy due to frequent scheduler invocations.
Dynamicity of the tile-based watermark pipeline: One way to better utilize GPU resources is to increase the batch size in the tile-based detection pipeline. However, while larger batches improve throughput, they almost inevitably increase latency. As shown in Fig. 2, simply enlarging the batch size in the naive tile-based pipeline quickly drives throughput to saturation—after which it plateaus or even slightly declines—while latency rises sharply. Moreover, most computation is concentrated in the decoding stage, i.e., extracting the watermark from each tile, which overlaps poorly with other operations. Consequently, decoding emerges as the primary bottleneck, limiting the benefits of larger batch sizes. To further analyze pipeline behavior, we evaluated different batch sizes under fixed stream allocations across stages. For a batch size of 256, setting the stream counts for preprocessing, decoding, and RS correction to $(1,1,16)$ yields a $1.43\times$ latency reduction compared to the single-stream baseline. In contrast, applying the same configuration to a batch size of 16 results in a slowdown, achieving only $0.86\times$ due to kernel-launch overhead. This discrepancy demonstrates that an optimal configuration for one setting does not necessarily transfer to another. These findings highlight the need for a strategy that can adapt to varying tile and batch sizes. Designing such an approach presents two key challenges. First, tile size and batch size alone are insufficient to accurately predict the resource demands of individual stages. Second, simply tuning the number of streams cannot balance stage execution times, often leading to pipeline bubbles or additional launch overhead. An adaptive, resource-aware method is therefore required to balance these competing factors and more effectively exploit GPU resources.
4 QRMark Algorithm
In this section, we first introduce how to pretrain a decoder to recover message bits from a specific size image tiles, and how an existing latent diffusion model decoder is fine-tuned to embed messages into tiles. We then explain how the pretrained decoder is used for message extraction, followed by using Reed-Solomon correction for accuracy.
4.1 Pre-training the tile-based watermark extractor
The pre-training process is consistent with the approach adopted in Stable Signature [9], with two key modifications: (1) tile-specific adaptations to enable training of a tile-based decoder, and (2) adjustments to the loss function to accommodate Reed–Solomon correction.
In Fig. 3(a), the pre-training pipeline consists of a convolutional encoder and a convolutional extractor. The encoder learns to embed imperceptible perturbations into the cover image, while the extractor is simultaneously trained to recover the embedded message from the watermarked image. Both encoder and decoder are optimized by the HiDDeN [52].
@@ 4.1 Pre-training the tile-based watermark extractor
Figure 3: Overview of QRMark Algorithm. (a) We begin by pre-training a watermarking encoder $H_{E}$ and an extractor $H_{D}$ to embed and extract a fixed bit sequence $m$. (b) Next, we fine-tune LDM’s decoder $D_{m}$ of the LDM auto-encoder with a Reed–Solomon encoded bit sequence $m_{s}$. (c) Finally, we apply the extractor $H_{D}$ to extract the hidden watermark message from the tiles of generated images.

### v2

@@ Abstract
Efficient and reliable detection of generated images is critical for the responsible deployment of generative models. Existing approaches primarily focus on improving detection accuracy and robustness under various image transformations and adversarial manipulations, yet they largely overlook the efficiency challenges of watermark detection across large-scale image collections. To address this gap, we propose QRMark, an efficient and adaptive end-to-end method for detecting embedded image watermarks. The core idea of QRMark is to combine QR Code–inspired error correction with tailored tiling techniques to improve detection efficiency while preserving accuracy and robustness. At the algorithmic level, QRMark employs a Reed–Solomon error correction mechanism to mitigate the accuracy degradation introduced by tiling. At the system level, QRMark implements a resource-aware multi-channel horizontal fusion policy that adaptively assigns more streams to GPU-intensive stages of the detection pipeline. It further employs a tile-based workload interleaving strategy to overlap data-loading overhead with computation and schedules kernels across stages to maximize efficiency. End-to-end evaluations show that QRMark achieves an average $2.43\times$ inference speedup over the sequential baseline.
@@ 1 Introduction
Recent advances in diffusion-based generative models [rombach2022high, ho2020denoising, nichol2021improved, song2020denoising] have progressively narrowed the gap between synthetic outputs and real images, making them increasingly difficult to distinguish. In this context, watermarking has become a crucial tool for verifying AI-generated content and supporting copyright attribution. Today, major social platforms handle billions of new visual uploads each day. For example, YouTube reports over 500 hours of video uploaded per minute on its official site [YouTubeHowItWorksSearch], and a recent measurement study estimates approximately 117 million TikTok posts per day [Steel2025TikTokHour]. Historical records also show Facebook receiving more than 350 million new photos per day [Wang2015FaceSearchAtScale]. The growing scale of online content creation highlights the need for watermark detection pipelines that not only achieve high accuracy, but also maintain robustness against adversarial perturbations and provide high efficiency when processing massive volumes of images.
@@
Among existing watermarking techniques, in-generation watermarking [fernandez2023stable, wen2023tree, wang2025sleepermark], where the watermark is incorporated directly into the generative process, has received the most attention and has emerged as the most widely adopted solution due to its accuracy and robustness. However, existing efforts largely overlook two critical practical factors of this method: the latency of watermark detection and the throughput required to reliably handle the massive volume of incoming images at scale. For example, even for relatively small $256\times 256$ images, the representative Stable Signature model requires about $30$ minutes to process $1$ million images—far from sufficient for social platforms that must process continuous streams of large-scale image and video content.
@@
One potential strategy to reduce computational overhead is tiling, where the detection process operates on a tile that still contains the embedded message. While tiling can substantially lower the computational cost of watermark detection, naive tiling does not necessarily yield proportional performance gains and often results in a noticeable drop in accuracy and robustness. Our experiments show that using the original pretrained Stable Signature decoder for tile-level detection achieves only $87.5\%$ bit accuracy, compared to $99.7\%$ with the full Stable Signature pipeline. Additional experiments consistently confirm that tiling inevitably degrades detection performance, underscoring the intrinsic trade-off between computational efficiency and accuracy. In this context, we require a method that combines low computational overhead with strong error-correction capability to offset the accuracy loss induced by tile-based watermarking. A familiar example is the QR Code, which remains reliably decodable even under partial occlusion or physical damage, thereby ensuring preservation of the stored data. Motivated by this observation, we examined the underlying mechanism and found that this resilience is largely enabled by Reed–Solomon (RS) error correction [reed1960polynomial, lim1978decoding]. We observe that RS correction holds potential for mitigating the accuracy and robustness degradation caused by tiling, as it can flexibly recover the original message bit string from partially corrupted outputs.
@@
Nevertheless, fully harnessing this potential introduces several system-level challenges. Our experiments show that applying a straightforward tiling strategy to Stable Signature yields modest performance gains ($1.17\times$ speedup) but still falls short of expectations. Moreover, RS correction is executed on the CPU, a common design choice since the algorithm is traditionally CPU-bound due to its many interdependent instruction flows (e.g., if–else branches), and thus benefits little from GPU acceleration, which relies on massively parallel execution. In a basic pipeline, this CPU stage can easily become a bottleneck, as it must synchronize with the rest of the detection process running on the GPU. These observations underscore the need to carefully redesign the pipeline to reduce computational overhead and ensure scalability under high-throughput workloads.
@@
Beyond algorithmic design, QRMark introduces several system-level innovations to maximize GPU utilization and reduce end-to-end latency. First, we implement an adaptive multi-channel horizontal fusion strategy that dynamically adjusts to image resolution and message length, ensuring balanced parallelism without incurring excessive kernel-launch overhead. Second, we design a resource-aware scheduling algorithm for tile-level workloads that distributes computation across GPU streams to mitigate imbalance and improve throughput. Finally, QRMark implements workload interleaving to overlap data loading with GPU execution, alleviating pipeline stall bubbles and enabling more efficient kernel orchestration.
@@
  • We design an end-to-end tile-based watermark detection framework that builds on this method while also incorporating system-level optimizations, including adaptive multi-channel horizontal fusion(§5) and resource-aware scheduling (§6), to effectively overcome the inefficiencies inherent in naive tile-based detection pipelines.
@@ 2 Background
Diffusion Watermarking: Diffusion generative models [dhariwal2021diffusion, ho2020denoising, nichol2021improved, song2020denoising] have renewed interest in watermarking as a means of ensuring content provenance. In this context, watermarking refers to embedding hidden identifiers into AI-generated images, enabling origin verification without relying on visible tags. While traditional methods embed secret pixel patterns [begum2020digital, wadhera2022comprehensive], modern approaches integrate watermarking directly into the generative process to improve robustness. Existing image watermarking techniques can be broadly classified into three major categories: post-generation, pre-generation, and in-generation. (1) Post-generation watermarking, where the watermark is embedded after the image has been generated [rahman2013dwt, zhang2019robust], is straightforward to implement but highly vulnerable in white-box scenarios, as the watermark can be easily removed. (2) Pre-generation watermarking, which embeds watermarks into all training images [yu2021artificial, zhao2023recipe], incurs prohibitively high computational costs for large-scale diffusion models. (3) In-generation watermarking, where the watermark is embedded during the image generation process [rahman2013dwt, tancik2020stegastamp, zhu2018hidden, luo2020distortion, al2007combined, jing2021hinet, zhang2020udh], distributes the signal across the entire image, making it difficult to remove without causing noticeable degradation. We adopt in-generation watermarking as the focus of our study, since it is widely accepted and provides superior robustness compared to alternative approaches.
System Support for Diffusion Watermarking: Prior system research on diffusion generation has primarily focused on accelerating the long sequential denoising pipeline, introducing optimizations such as kernel fusion, pipeline parallelism, stream batching, and patch parallelism [tensorrt, li2024distrifusion, kodaira2023streamdiffusion, zhong2023brief]. These methods improve forward sampling efficiency but assume a monolithic, step-driven workload with strong inter-step dependencies, leaving decode-time scheduling demands unaddressed. Interleaving across CPU and GPU stages is crucial for diffusion watermark detection to achieve lower latency, yet this aspect is absent from prior diffusion system designs. System-level efforts in watermarking, on the other hand, have largely targeted post-hoc embedding and detection, typically modeled as simple feed-forward computations with regular GPU kernels and minimal CPU involvement [hussain2022faststamp, zhang2025robust, uchida2017embedding]. These methods assume image-level operators and homogeneous kernels, making them well-suited for global transforms but incompatible with patch-based workloads. In contrast, diffusion watermark detection relies on patch extraction, fragmenting computation into numerous small kernels and inducing irregular device–host interactions. Therefore, diffusion watermark detection presents new opportunities when treated as an integrated workload rather than two disjoint problems.
@@ 3 Motivation
Existing watermarking systems [fernandez2023stable, wen2023tree, feng2024aqualora, wang2025sleepermark, zhao2023recipe, min2024watermark] largely overlook efficiency considerations, leaving substantial opportunities for further computational optimization. To better understand how watermarking systems can be made more efficient, we focus on a representative model, Stable Signature [fernandez2023stable], and systematically explore both algorithmic and system-level optimizations carefully built on top of it.
Divisible image watermark detection with tiling: By analyzing the Stable Signature watermark detection pipeline, we identify an opportunity to exploit tiling. Since an image is inherently divisible, it can be partitioned into tiles for localized processing. Empirical evidence from Stable Signature shows that under a cropping attack where only $10\%$ of the original area is preserved, the model still achieves over $95\%$ detection accuracy [fernandez2023stable]. This observation suggests that, even without explicit tile-level training, the message extractor is capable of recovering the embedded message from partial image regions. Leveraging this property, we introduce tiling into the detection process. Tiling is a well-established technique in system optimization [xu2009tiling, rolih2024divide, jangda2020model, ozge2019power], widely used to reduce both computational and memory overhead. In the context of image watermark detection, our experiment (Fig. 2(a)) demonstrates that integrating tiling into Stable Signature yields notable scalability benefits: with tiling, the system achieves up to a $1.23\times$ speedup over the non-tiled baseline, whereas the non-tiled baseline achieves at most $1.06\times$ across varying batch sizes. This result highlights the superior scalability brought by tiling. However, tiling inevitably introduces an efficiency–accuracy trade-off, as partitioning the image may compromise detection fidelity. To address this challenge, we draw inspiration from the design of QR Codes, where Reed–Solomon error correction ensures reliable decoding under data loss and corruption. This motivates our design of QRMark, which integrates Reed–Solomon correction to offset the accuracy degradation caused by tiling and achieve a more favorable balance between efficiency and robustness.
@@
Performance dilemma of watermark image tiling: Although replacing full-image processing with tile-level extraction is expected to reduce per-image overhead and improve detection efficiency, our experiments on 1,000 images from the MS-COCO [lin2015microsoftcococommonobjects] dataset confirm that the naive tile-based design brings only limited improvements: with a tile size of 64 and a batch size of 16, the pipeline achieves a $1.17\times$ speedup over the non-tiled baseline. However, this gain falls short of expectations, prompting the question of what factors constrain the potential of the tile-based approach. To answer this, we conducted profiling experiments with a default tile size of 64, a message length of 48, and a batch size of 16. The results reveal two primary bottlenecks. First, GPU utilization remains low: Stable Signature leverages only $31.6\%$ of available GPU capacity. Second, the Reed–Solomon correction, executed on the CPU, introduces additional overhead and occasionally stalls the entire pipeline. Further experiments also show that the transform phase preceding detection in the original Stable Signature implementation launches a large number of fine-grained kernels. These fragmented kernels originate from preprocessing operations such as cropping, resizing, and normalization. The fragmented design incurs substantial kernel-launch overhead and prevents the GPU from achieving high occupancy due to frequent scheduler invocations.
@@
Figure 3: Overview of QRMark Algorithm. (a) We pretrain a watermarking encoder $H_{E}$ and an extractor $H_{D}$ with customized tiling settings to embed and extract a fixed bit sequence $m$, the encoder $H_{E}$ is then discard; (b) We fine-tune latent diffusion model’s decoder $D_{m}$ with a Reed–Solomon encoded bit sequence $m_{s}$; (c) We apply the extractor $H_{D}$ to extract the hidden watermark message from the tiles of generated images. A more detailed breakdown on the detection pipeline is shown in Fig. 4.
Dynamicity of the tile-based watermark pipeline: One way to better utilize GPU resources is to increase the batch size in the tile-based detection pipeline. However, while larger batches improve throughput, they almost inevitably increase latency. Fig. 2 shows that simply enlarging the batch size in the naive tile-based pipeline quickly drives throughput to saturation, after which it plateaus or even slightly declines, while latency rises sharply. Moreover, most computation is concentrated in the decoding stage, i.e., extracting the watermark from each tile, which overlaps poorly with other operations. Thus, decoding emerges as the primary bottleneck, limiting the benefits of larger batch sizes. To further analyze pipeline behavior, we evaluated different batch sizes under a fixed multi-channel horizontal fusion strategy across stages. For a batch size of 256, setting the stream counts for preprocessing, decoding, and RS correction to $(1,1,16)$ yields a $1.43\times$ latency reduction compared to the single-stream baseline. In contrast, applying the same configuration to a batch size of 16 results in a slowdown, achieving only $0.86\times$ due to kernel-launch overhead. This discrepancy demonstrates that an optimal configuration for one setting does not necessarily transfer to another. These findings highlight the need for a strategy that can adapt to varying tile and batch sizes. Designing such an approach presents two key challenges. First, tile size and batch size alone are insufficient to accurately predict the resource demands of individual stages. Second, simply tuning the number of streams cannot balance stage execution times, often leading to pipeline bubbles or additional launch overhead. An adaptive, resource-aware method is therefore required to balance these competing factors and more effectively exploit GPU.
4 Tiled Watermark Detection Optimization
While the proposed approach in Stable Signature [fernandez2023stable] achieves high accuracy in watermark recovery, the deployment of the detection pipeline introduces system-level constraints. First, the watermark detection process is compute-intensive: the execution of the message decoder $H_{D}$ is dominated by kernels that operate on dense pixel data from each watermarked tile. Second, the computational cost is intrinsically tied to the spatial resolution of the inputs. In text-to-image diffusion pipelines, it is standard practice to partition high-resolution images into fixed-size tiles so that watermark encoders and decoders can operate locally on spatial regions[wang2025timestep, ding2024patched, zhang2023sine, madar2025tiled, sun2025patchedserve]. We therefore adapt a similar technique to the watermark detection pipeline. Each time, we only need to use a single tile of the watermarked image instead of the whole image. Under this tiling scheme, the tile size directly determines the workload of these kernels, and increasing the tile dimensions to improve robustness leads to a proportional increase in FLOPs and decoding latency.
Motivated by these constraints, we next describe how QRMark is designed to support efficient and scalable watermark extraction. In this section, we first introduce the pretraining of a decoder to recover message bits from fixed-size image tiles and the fine-tuning of an existing latent diffusion model decoder to embed messages into these tiles. We then present the algorithmic workflow of our message extraction pipeline and conclude by explaining how Reed–Solomon correction is applied to further improve accuracy.
4.1 Pre-training Tiled Watermark Extractor
The pre-training process is consistent with the approach adopted in Stable Signature [fernandez2023stable], with two key modifications: (1) tile-specific adaptations to enable training of a tile-based decoder, and (2) adjustments to the loss function to accommodate Reed–Solomon correction.